ratio question:
 
Divide Rs 1200 among A,B,C so tha A gets 2/3 of what B gets and B gets1/4 of what C gets. Find the share of each

Grade:

1 Answers

srikar
9 years ago
A + B + C = 1200
A = 2A/3         B = C/4     A= C/6
B= 3A/2     C=6A
==> A + B + C = A+ 3A/2 + 6A = 1200 ==>17A/2 =1200
A= 2400/17Rs     B= 3600/17Rs      C = 14400/17Rs